### v1

# MolErr2Fix: Benchmarking LLM Trustworthiness in Chemistry via Modular Error Detection, Localization, Explanation, and Revision

Abstract
Large Language Models (LLMs) have shown growing potential in molecular sciences, but they often produce chemically inaccurate descriptions and struggle to recognize or justify potential errors. This raises important concerns about their robustness and reliability in scientific applications. To support more rigorous evaluation of LLMs in chemical reasoning, we present the MolErr2Fix benchmark, designed to assess LLMs on error detection and correction in molecular descriptions. Unlike existing benchmarks focused on molecule-to-text generation or property prediction, MolErr2Fix emphasizes fine-grained chemical understanding. It tasks LLMs with identifying, localizing, explaining, and revising potential structural and semantic errors in molecular descriptions. Specifically, MolErr2Fix consists of 1,193 fine-grained annotated error instances. Each instance contains quadruple annotations, i.e,. (error type, span location, the explanation, and the correction). These tasks are intended to reflect the types of reasoning and verification required in real-world chemical communication. Evaluations of current state-of-the-art LLMs reveal notable performance gaps, underscoring the need for more robust chemical reasoning capabilities. MolErr2Fix provides a focused benchmark for evaluating such capabilities and aims to support progress toward more reliable and chemically informed language models. All annotations and an accompanying evaluation API will be publicly released to facilitate future research.
MolErr2Fix: Benchmarking LLM Trustworthiness in Chemistry via Modular Error Detection, Localization, Explanation, and Revision
Yuyang Wu1††thanks: Equal contribution, Jinhui Ye1,*, Shuhao Zhang1,*, Lu Dai2, Yonatan Bisk1††thanks: Corresponding authors, Olexandr Isayev1, 1 Carnegie Mellon University 2Hong Kong University of Science and Technology {yuyangwu, jinhuiye, shuhaozh}@andrew.cmu.edu, ldaiae@connect.ust.hk, ybisk@cs.cmu.edu, olexandr@cmu.edu
1 Introduction
Large Language Models (LLMs) have achieved remarkable success in natural language tasks, but recent studies highlight significant limitations when these models are applied to complex chemistry problems Jablonka et al. (2022); Hatakeyama-Sato et al. (2023); Bran et al. (2025). Even state-of-the-art LLMs struggle with domain-specific knowledge and often produce misleading or incorrect outputs in molecular science applications Guo et al. (2023). One prominent issue is the hallucination of incorrect details when LLMs describe molecules (“molecular captioning”). In particular, models fine-tuned to translate between chemical structures and text frequently generate factually incorrect or nonsensical information about molecular structures or their properties Lu et al. (2024). As illustrated in Figure 1, LLMs often produce fluent and grammatically correct molecular captions, yet misinterpret the underlying SMILES Weininger (1988) strings—for example, failing to recognize the correct number of atoms or functional groups. Such chemical errors are substantially more critical than surface-level linguistic inaccuracies. More alarmingly, most LLM models consistently fail to recognize their own factual or logical errors in a given translation result Kamoi et al. (2024).
Previous approaches to evaluate the chemical understanding of models have predominantly relied on indirect assessments through downstream tasks such as molecular property prediction or molecular question answering Lu et al. (2024). This reliance on indirect evaluation was, in part, a consequence of the limited generative capabilities of earlier models and a scarcity of sophisticated tools for effectively assessing the quality of generated molecular text: traditional text generation metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) prioritized linguistic overlap over restricted outputs, obscuring whether a model genuinely reasons over molecular semantics. Recently, with the significant advancements in LLM generation, efforts have emerged to directly probe their understanding by having them produce textual descriptions of molecules (e.g., SMILEStext). Platforms like MolCap-ArenaEdwards et al. (2024), for example, facilitate the comparison of different models’ descriptive capabilities through ranking methodologies such as pairwise comparisons, but still lack the granularity to provide deep insights into the specific nature of an LLM’s misunderstandings of molecular features. They typically do not pinpoint whether a model fails to identify functional groups, miscounts atoms, or makes other critical chemical errors, thus offering limited actionable feedback for targeted model improvement.
In real-world scenarios, LLMs are expected to serve as collaborative chemical assistants—working with scientists, generating hypotheses, and accelerating discovery Bran et al. (2025). Fulfilling this role requires more than fluent molecule descriptions or one-shot answers; models must also detect, localize, explain, and revise their own chemical errors. However, no existing benchmark reveals whether an LLM can self-verify or correct flawed chemical claims—an ability that underpins any trustworthy scientific assistant.
To address this gap, we introduce MolErr2Fix, a structured benchmark that systematically evaluates LLMs’ capabilities in chemical error detecting and reasoning. Given an initial molecular description, LLMs are tasked with: 1) detecting whether errors exist, 2) localizing the erroneous span within the text, 3) explaining the violated chemical principle, and 4) generating a corrected, chemically valid revision. Unlike existing evaluations focused on fluency or factual recall, MolErr2Fix probes deeper reasoning and diagnostic abilities. It offers fine-grained insights into whether LLMs can leverage domain knowledge to identify and correct errors in molecular descriptions.
By leveraging this comprehensive benchmark, our evaluation reveals critical shortcomings in current models (including GPT-4o and chemistry-specific ChemLLM Zhang et al. (2024)). While a few models occasionally succeed at detecting errors, they frequently fail to localize, explain, or correct them accurately. These findings underscore that even the most advanced LLMs lack some mechanisms for chemical error discovery and recovery.
Our main contributions can be summarized as follows.
We propose MolErr2Fix, the first benchmark explicitly designed to evaluate LLMs’ ability to detect and recover from chemical errors in molecular descriptions. Unlike prior tasks focused solely on generation or classification, our benchmark directly assesses models’ diagnostic reasoning and reliability in a chemically grounded setting.
MolErr2Fix introduces a structured chain-like, four-stage Error-to-Fix evaluation pipeline—detection, localization, explanation, and revision—offering a fine-grained and progressive framework for assessing LLMs in chemical reasoning.
We provide detailed error analysis and propose evaluation protocols to benchmark and progress toward chemically reliable LLMs.
2 Related Works
2.1 Molecule Captioning and Translation
Early work on bridging molecules with natural language established tasks like molecule captioning and cross-modal retrieval, with Edwards et al. (2021) introducing ChEBI-20 and Text2Mol for molecule-text alignment. Various cross-modal models (Su et al., 2022; Christofidellis et al., 2023; Luo et al., 2023; Liu et al., 2023b, c; Zhang et al., 2024) bridged molecular and natural language via seq2seq tasks (e.g., molecule captioning, de novo generation) and contrastive tasks (e.g., cross-modal retrieval). Generative models like MolT5 (Edwards et al., 2022) and BioT5 (Pei et al., 2023), and contrastive models like MoMu (Su et al., 2022) and MoleculeSTM (Liu et al., 2023a), advanced these tasks. MolReGPT (Li et al., 2025) enabled in-context learning with GPT-3.5/4. However, these works focus on generation or retrieval, not chemical correctness.
2.2 Large Language Models for Chemistry
Studies on LLMs’ chemical understanding have examined tasks like retrosynthesis planning and functional group identification. Bran et al. (2025) integrated GPT-4 with algorithmic search to enable retrosynthesis planning, demonstrating advanced synthetic route analysis. However, Yik and Dood (2024) found LLMs frequently generate incorrect chemical outputs despite fluency. Similarly, Li et al. (2024) identified factual errors in MolT5-generated descriptions, undetected by surface-level metrics. MoleculeQA (Lu et al., 2024), with over 60,000 questions, showed that specialized models achieve < 50% accuracy in factual comprehension, underscoring persistent challenges to ensure chemical reliability.
Recent LLMs for chemistry include benchmark studies (White et al., 2023; Guo et al., 2023; Jablonka et al., 2024) evaluating GPT-4 and LLaMA, Guo et al. (2023) noting chemically implausible outputs. Fine-tuning datasets like Mol-Instructions (Fang et al., 2023), Drugchat (Liang et al., 2023), and MolOpt-Instructions (Ye et al., 2025) aim to improve LLMs, with limited gains. SMolInstruct (Yu et al., 2024) enables LlaSMol that surpasses state-of-the-art LLMs. Nevertheless, no prior benchmark directly targets the detection and correction of chemical errors in natural language statements. Our MolErr2Fix fills this gap by challenging LLMs to identify and fix molecular misstatements, thereby exposing the remaining barriers to trustworthy chemical reasoning.
3 The MolErr2Fix Benchmark
We introduce MolErr2Fix, a comprehensive molecular benchmark that systematically evaluates Large Language Models (LLMs) by assessing their deep understanding of chemical knowledge through an innovative paradigm of fine-grained error detection, localization, explanation, and correction tasks, which traditionally rely on chemist-level expertise.
3.1 Task Formulation
The MolErr2Fix benchmark evaluates models across four integral tasks, each targeting specific dimensions of chemical reasoning and knowledge. This facilitates a comprehensive assessment of model performance from error identification through to correction. These tasks challenge models in a structured manner, with Figures 2 illustrating details of the four tasks.
Error Detection evaluates whether the model can identify the presence of chemical errors within a given flawed description () of a molecular structure (SMILES). The model is expected to output a predicted set of error types that it believes are present, chosen from the six predefined error categories (). Performance is measured using standard multi-label classification metrics: Precision, defined as , and Recall, defined as , where is the set of ground-truth error types. The harmonic mean F1 Score is reported to balance these two aspects.
Error Localization measures the model’s ability to locate the specific erroneous text spans in the flawed description () for a given molecular structure. The model should output a set of predicted error spans , where each denotes a contiguous text segment in predicted to be erroneous. We evaluate the span-level overlap between these predicted spans and the gold-standard spans using Intersection-over-Union (IoU). Specifically, we assess Recall@IoU (the proportion of gold spans for which a predicted span achieves an , with set at 0.5 and 0.7) and the Average IoU, calculated as .
Error Explanation assesses whether the model can explain why a given gold error span within a flawed description of a molecule (SMILES) is chemically incorrect. The model is required to generate a concise natural language explanation (typically 1–2 sentences). The quality of these explanations is quantitatively evaluated by comparing them against expert-written references using GPT-4o as an automatic evaluator for semantic equivalence. This is reported as a Match Rate: , where is an indicator function signifying semantic equivalence as judged by GPT-4o.
Error Correction tests the model’s ability to generate accurate textual corrections for an identified gold error span in a flawed description corresponding to a given molecular structure. The model’s output is a corrected version that replaces with a chemically accurate alternative. These corrections are evaluated for textual similarity against expert-validated revisions using BLEU scores. Furthermore, a Correction Score, based on GPT-4o judging semantic equivalence (), is used to assess deeper accuracy.
3.2 Dataset Curation
The construction of MolErr2Fix follows a two–stage workflow: (i) data collection & preparation and (ii) expert annotation & verification, as depicted in Figure 2, with complete prompt templates and decoding parameters provided in the appendix.
Data Collection and Preparation. We begin with the ChEBI-20 dataset and select molecules containing fewer than 100 atoms. This threshold balances molecular complexity and the input–length limits of contemporary sequence-to-sequence generators.
These molecules, represented in Simplified Molecular Input Line Entry System (SMILES) format, were then used to generate captions via LLMs including GPT-4o, o3-mini, Claude-Sonnet 3.7, Gemini 1.5, and the chemistry-specialised MolT5-high, under a zero-shot prompt adapted from MolReGPT. The resulting free-form descriptions are fluent but frequently contain chemically incorrect spans; we therefore collect them as Problematic Captions to be diagnosed in stage (ii).
Meanwhile, to establish a structured annotation scheme, a panel of chemistry experts mapped LLM failure modes onto IUPAC nomenclature and CAS indexing rules, iteratively consulting PubChem until consensus was reached. The resulting six error types for molecular captioning are: Functional Group/Substituent Errors (misidentifying molecular substructures or functional groups), Classification Errors (incorrectly categorizing the chemical species), Derivation Errors (erroneously linking to chemical precursors or derivatives), Stereochemistry Errors (inaccurately assigning stereochemical configurations like R/S or E/Z), Sequence/Composition Errors (miscounting molecular components or chain lengths), and Indexing Errors (incorrectly locating substituents or features). These error types are not only grounded in established nomenclature practice, but are also designed to comprehensively assess an LLM’s depth of understanding across numerical, structural, and semantic aspects of chemical information. This preparatory step laid the groundwork for the subsequent expert annotation stage, where problematic captions would be systematically analyzed and corrected. More detailed definitions and examples of the molecule captioning errors are in the Appendix.
Expert Annotation and Verification. The second stage involved meticulous annotation and verification by chemistry experts. Guided by predefined annotation criteria and workflows, these experts systematically analyzed the LLM-generated captions against the corresponding molecular structures. For each identified discrepancy, their annotation pipeline comprised of the four key tasks.
The initial step, Error Localization, involved experts precisely pinpointing the specific text spans within the caption that contained an error. Following this, the Error Classification task required them to identify the type of chemical error based on the predefined six-category taxonomy: functional group errors, classification errors, derivation errors, stereochemistry inaccuracies, sequence inconsistencies, and indexing errors. Subsequently, for Error Explanation, the experts articulated a concise and chemically precise explanation for why the identified span was incorrect, referencing relevant chemical principles or structural features. The final task, Error Correction, involved providing a corrected version of the erroneous text span or, if necessary, the entire description, ensuring chemical accuracy and consistency with the reference molecule.
To ensure the validity of their assessments and corrections, the experts consistently utilized a suite of standard reference tools, including molecular visualization software (e.g., RDKit), authoritative chemical databases (such as PubChem), and relevant research articles. Furthermore, a rigorous verification phase, incorporating a cross-check and recheck protocol among at least two experts for each instance, was implemented to maintain high inter-expert agreement and overall annotation quality. This expert-driven annotation and verification phase was crucial in transforming the raw, potentially flawed LLM outputs into a high-quality benchmark dataset. At the end, each instance includes the molecule, the erroneous description, and detailed, verified annotations regarding the error’s type, location, explanation, and correction.
3.3 Dataset Statistics
| Statistics | Number |
| Total Molecules | 525 |
| Total Annotated Errors | 1193 |
| Error Types | 6 |
| Models Evaluated | 5 |
| Error Type Distribution | |
| Functional Group/Substituent Errors | 318 (26.6%) |
| Classification Errors | 303 (25.4%) |
| Derivation Errors | 220 (18.4%) |
| Sequence/Composition Errors | 152 (12.7%) |
| Stereochemistry Errors | 111 (9.3%) |
| Indexing Errors | 89 (7.5%) |
| Error Frequency | |
| Peak Errors per Molecule | 3 (225 instances) |
| Average Errors per Molecule | 2.27 |
| Textual Component Length (characters) | |
| Average Description Length | 652.2 |
| Average Error Segment Length | 419.8 |
| Average Explanation Length | 167.6 |
| Average Correction Length | 211.0 |
The MolErr2Fix benchmark is constructed from 525 unique molecules, leading to 1,193 meticulously annotated errors, categorized into six distinct types, as shown in Table 1. The distribution of these errors highlights specific challenges for Large Language Models (LLMs) in the chemical domain. Functional Group/Substituent Errors are the most prevalent, accounting for of all identified issues, closely followed by Classification Errors at . This indicates that LLMs often struggle with accurately identifying and describing fundamental chemical moieties and correctly categorizing molecules. Other error types include Derivation Errors (), Sequence/Composition Errors (), Stereochemistry Errors (), and Indexing Errors (). The textual characteristics of the benchmark components also provide insight: average description lengths are characters, while the average length of an identified erroneous segment is characters. This considerable length for error segments implies that errors are often complex and deeply embedded within the surrounding text. Expert-provided reasoning explanations average characters and corrections average characters, underscoring the detailed nature of the annotations and the effort required for rectification, reflecting the diverse complexity in error justification and correction of our benchmark.
| Models | ErrorDetection | ErrorLocalization | ErrorExplanation | ErrorRevision | ||||||
| Recall | Precision | F1 | IoU≥0.5 | IoU≥0.7 | IoUavg | BLEU | GPTscore | BLEU | GPTscore | |
| 1) LLM Models | ||||||||||
| gpt-4 | 9.7 | 48.3 | 16.1 | 21.8 | 15.3 | 15.7 | 1.4 | 12.8 | 9.6 | 3.5 |
| gpt-4o | 52.3 | 49.8 | 51.0 | 35.4 | 20.9 | 23.3 | 1.6 | 13.7 | 4.7 | 1.7 |
| qwenvl-72B | 45.6 | 47.4 | 46.5 | 47.2 | 29.1 | 31.7 | 0.8 | 3.2 | 1.4 | 0.2 |
| gemini-2.0-flash | 36.2 | 54.4 | 43.5 | 27.9 | 16.1 | 18.2 | 1.1 | 10.0 | 1.5 | 12.5 |
| deepseek-chat | 35.0 | 58.3 | 43.8 | 50.2 | 34.1 | 35.8 | 1.2 | 15.2 | 0.9 | 4.5 |
| Claude-3.7-sonnet | 51.9 | 53.3 | 52.6 | 31.6 | 12.8 | 16.1 | 1.9 | 20.9 | 0.5 | 7.9 |
| Grok-3 | 76.8 | 58.7 | 66.6 | 41.1 | 25.6 | 26.7 | 1.9 | 21.6 | 0.4 | 7.4 |
| 2) LLM Reasoner Models | ||||||||||
| o3-mini | 29.8 | 56.8 | 39.1 | 29.0 | 14.5 | 17.4 | 0.8 | 41.1 | 1.8 | 2.8 |
| o4-mini | 44.8 | 53.9 | 53.5 | 24.1 | 12.7 | 14.9 | 0.4 | 43.1 | 1.6 | 3.6 |
| deepseek-reasoner | 52.8 | 54.3 | 48.9 | 31.2 | 17.3 | 20.1 | 0.4 | 18.4 | 1.8 | 2.0 |
| 3) Domain-specific LLM Models | ||||||||||
| LlaSMol | 34.2 | 3.4 | 6.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| ChemLLM-chat | 40.3 | 4.2 | 7.5 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| ChemLLM-chat-DPO | 100.0 | 4.2 | 8.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| 3) LLM Models + 5-shot Prompt | ||||||||||
| gpt-4 | 25.7 | 51.5 | 34.3 | 2.7 | 2.5 | 2.5 | 1.6 | 8.8 | 11.1 | 4.5 |
| gpt-4o | 50.7 | 53.9 | 52.3 | 60.1 | 47.4 | 49.1 | 1.6 | 12.8 | 5.7 | 2.1 |
| o3-mini | 36.5 | 60.5 | 45.6 | 38.5 | 25.3 | 27.4 | 0.9 | 20.0 | 2.7 | 5.0 |
| o4-mini | 50.3 | 52.4 | 51.3 | 41.1 | 28.6 | 30.6 | 0.5 | 24.3 | 2.5 | 4.6 |
| qwen-plus | 29.0 | 46.4 | 35.7 | 39.7 | 29.7 | 31.1 | 1.3 | 5.1 | 3.2 | 4.8 |
| gemini-2.0-flash | 40.0 | 47.1 | 43.2 | 55.1 | 45.7 | 47.1 | 1.7 | 9.9 | 6.1 | 1.8 |
4 Experiments
The main purpose of baseline experiments is to investigate LLMs’ performance in detecting, localizing, explaining, and revising errors in molecular descriptions. The evaluation encompasses four tasks: Error Detection, Error Localization, Error Explanation, and Error Revision.
4.1 Baselines
We categorize the baseline models into three groups based on their architectural design and adaptation strategies:
General-purpose LLMs: This group includes gpt-4 Achiam et al. (2023), gpt-4o Hurst et al. (2024), qwen1-72B Bai et al. (2023), gemini-2.0-flash Team et al. (2023), deepseek-chat Liu et al. (2024), Claude-3.7-sonnet Anthropic (2024), and Grok-3. These models, developed for broad natural language understanding tasks, serve as a foundation for assessing general error-handling abilities in molecular contexts.
Reasoning-enhanced LLMs: Represented by o3-mini, o4-mini, and deepseek-reasoner Guo et al. (2025), these models are either designed or fine-tuned to excel in reasoning tasks. They serves to test the hypothesis that enhanced logical inference capabilities improve performance on chemically intricate error-handling tasks.
Domain-specific LLMs: This category features models that have been specifically adapted for the chemical domain. We include ChemLLM-chat ChemLLM-chat-DPO Zhang et al. (2024) and LlaSMol Yu et al. (2024) to allow for an assessment of the influence brought by specialized domain knowledge in addressing molecular errors.
Few-shot learning variants: We evaluate a subset of models—-gpt-4, gpt-4o, gemini-2.0-flash, o3-mini, o4-mini, and qwen-plus—in a few-shot setting, providing five task-specific examples for each. This category aims to examine the impact of in-context learning on task performance.
These categories enable a comprehensive comparison across model types and adaptation methods. We evaluate the four tasks of the MolErr2Fix benchmark as defined in Section 3.1. For Error Detection, we use Precision, Recall, and F1 Score. For Error Localization, we report Recall@IoU, Recall@IoU, and Average IoU. For Error Explanation and Error Revision, we apply BLEU and GPTScore to measure textual and semantic accuracy, respectively. All models are evaluated in a zero-shot setting, except for the few-shot learning variants. Details about the prompting strategies are provided in the Appendix.
4.2 Main Results
This section presents a comprehensive comparison of various LLMs on the MolErr2Fix benchmark, with detailed results presented in Table 2. Our key findings are summarized as follows:
Performance across tasks. Error Detection emerged as the most tractable task, where Grok-3 achieved the highest F1 Score of 66.6%, followed by Claude-3.7-sonnet (52.6%). In contrast, Error Revision proved to be the most challenging task, with no model surpassing a BLEU score of 11.1% (gpt-4, 5-shot). Error Localization and Error Explanation exhibited intermediate difficulty, with peak performances of 60.1% Recall@IoU (gpt-4o, 5-shot) for the former and a 43.1% GPT Score (o4-mini) for the latter, respectively. This observed hierarchy suggests an increasing level of difficulty for tasks that demand more profound chemical reasoning.
Limitations of domain-specific models. We assess three domain-specific 7B models—LlaSMol, ChemLLM-chat, and ChemLLM-chat-DPO—on our benchmark. These models are fine-tuned for narrow tasks like molecular captioning, and lack the dialogue or reasoning abilities required for multi-step instructions. As a result, they perform poorly across all tasks, especially in localization, explanation, and revision. The results highlight that domain alignment alone is insufficient without instruction-following capabilities.
Impact of Reasoning-enhanced. Reasoning-enhanced models, such as o4-mini (F1: 53.5% in Error Detection) and deepseek-reasoner, consistently outperformed general-purpose LLMs like gpt-4 (F1: 16.1%). This indicates that specialized training for logical inference significantly enhances error-handling capabilities within the molecular domain.
Impact of few-shot learning. Few-shot learning variants, notably gpt-4o (5-shot), demonstrated marked improvements, particularly in Error Localization, achieving IoU of 60.1% compared to 35.4% in a zero-shot setting. However, the performance gains were limited in Error Explanation and Error Revision, suggesting that in-context learning alone may be insufficient for these more complex generative tasks.
4.3 In-depth Analysis of Failure Modes
While the main results provide a quantitative overview of model performance, a more granular investigation is crucial for understanding the specific weaknesses and challenges in molecular captioning. This section therefore delves into the predominant failure modes, the models’ capabilities in error discovery and correction, and an exploration of potential underlying causes for the observed limitations.
4.3.1 Identifying Dominant Failure Modes
Figure 3 exposes three systemic weaknesses shared by all five models—Functional Group, Derivation, and Classification errors. Functional-group errors are the most pervasive, with GPT-4o and Claude mislabeling local substructures in of cases, underscoring the difficulty of binding graph-level patterns (e.g., sulfonate vs. sulfonamide) to precise lexemes. Derivation errors, frequently exceeding 70% on Gemini and GPT-4o, reveal an equally stark gap in reasoning over reaction or conjugation semantics: models rarely grasp that methyl salicylate is an ester of salicylic acid or that deprotonation conserves heavy-atom topology. Even MolT5—which is pre-trained on chemistry corpora—shows a non-trivial 24 % misclassification rate, indicating persistent difficulty coupling global topology with formal chemical taxonomies.
4.3.2 Limits of Error Discovery
The MolErr2Fix results in Figures 4–5 confirm that LLMs detect linguistically salient flaws but falter when numerical or spatial reasoning is required. Sequence errors (E5) are identified with a poor and localized with ; Indexing errors (E6) performs even worse (, ). These patterns highlight reliance on surface cues: Training data emphasize stylistic fluency, not atom counts or positional descriptors, leaving LLMs blind to subtle violations of IUPAC nomenclature or ring-index conventions.
4.3.3 Error Recovery Remains Elusive
Figure 6 benchmarks GPT-o4-mini—the best performer in our study—on explanation and revision. While semantic parity scores (GPT Score ) suggest it can paraphrase expert rationales, BLEU scores below 2 reveal brittle alignment with gold explanations. Revision is weaker still: BLEU and GPT Score for both E5 and E6 underscore that correcting numeric chains or ortho/meta/para indices remains largely out of reach.
4.3.4 Exploring Potential Underlying Causes
While the exact reasons for the low performance of LLMs in molecule captioning remain uncertain, we propose several possible explanations from a chemical perspective:
Certain essential domain knowledge, such as the indexing rules of atoms in fused-ring systems, is rarely detailed even in textbooks, let alone in the corpora typically used to train LLMs. This scarcity of specialized information makes it challenging to identify and address deficiencies in indexing tasks.
Stereochemistry information in SMILES representations is not explicitly indicated by specific symbols; instead, chiral configurations must be inferred from the relative order of fragment representations around an ’@’ symbol. Such inferential demands pose significant challenges for non-reasoning-based models in correctly distinguishing chiral structures.
Historically, many chemical species and fragments have multiple valid names, creating inherent ambiguities. Such ambiguities may increase the complexity for LLMs to consistently select appropriate terminology.
5 Conclusion and Future Work
MolErr2Fix introduces a four-stage audit—detection, localization, explanation, and correction—that forces LLMs to reason chemically rather than linguistically. Our evaluation reveals shortcomings in existing models: even state-of-the-art systems spot obvious flaws but falter when asked to locate, justify, or repair them, exposing a gap between fluent language generation and dependable chemical reasoning. We advocate (1) chemistry-centric pretraining architectures, (2) self-reflection loops for iterative debugging, and (3) an expanded benchmark covering richer chemistries and error modes. MolErr2Fix thus maps today’s limits and charts a path toward scientifically trustworthy LLMs.
Limitations
This study on language models’ chemical error detection has several limitations: the dataset only includes molecules under 100 atoms; six predefined error categories don’t capture all possible errors; and expert annotations may contain subjective biases. The evaluation is limited to certain model types, uses only a few-shot examples, and relies on metrics that may not fully capture chemical accuracy. Future research should expand molecular complexity coverage, error types, model diversity, and develop chemistry-specific evaluation metrics.